A MAN IS 3 TIMES AS OLD AS HIS SON,4 YEARS AGO,THE SUM OF THEIR AGES WAS 24.FIND THEIR PRESENT AGES
Answer by pj66(7)

Let x be the present age of son.
=> father's age = 3*x
Four years ago,
father's age = 3*x - 4 and son's age = x-4
Sum of their ages is 24
=> 3*x-4+x-4=24
=> 4*x-8=24
=> 4*x=32
=> x=8
Hence, father's present age is 24.